OpenStudy (anonymous):

I need help with conics. Major help actually...Quickly can someone help me get these conic equations into standard form?? Medal and Fan rewarded.

OpenStudy (anonymous):

show them

OpenStudy (anonymous):

One second :)

OpenStudy (anonymous):

-2y^2+x-4y+1=0 -25x^2+y^2-100x-125=0 4x^2+4y^2-20x-32y+81=0

OpenStudy (anonymous):

oh wow

OpenStudy (anonymous):

Do you not know how to do them??

OpenStudy (anonymous):

there is just alot hold on

OpenStudy (anonymous):

Yessir.

OpenStudy (amistre64):

standard form is really about completing squares in x and y